# Fine-grained Hallucination Detection and Editing for Language Models

### v3

@@ Abstract
Large language models (LMs) are prone to generate factual errors, which are often called hallucinations. In this paper, we introduce a comprehensive taxonomy of hallucinations and argue that hallucinations manifest in diverse forms, each requiring varying degrees of careful assessments to verify factuality. We propose a novel task of automatic fine-grained hallucination detection and construct a new evaluation benchmark, FavaBench, that includes about one thousand fine-grained human judgments on three LM outputs across various domains. Our analysis reveals that ChatGPT and Llama2-Chat (70B, 7B) exhibit diverse types of hallucinations in the majority of their outputs in information-seeking scenarios. We train Fava, a retrieval-augmented LM by carefully creating synthetic data to detect and correct fine-grained hallucinations. On our benchmark, our automatic and human evaluations show that Fava significantly outperforms ChatGPT and GPT-4 on fine-grained hallucination detection, and edits suggested by Fava improve the factuality of LM-generated text. [^1] Our code, data, and demo are available at https://fine-grained-hallucination.github.io/
Fine-grained Hallucination Detection and Editing for Language Models
[…]
@@ 1 Introduction
Large language models (LMs; Brown et al. 2020) can generate highly fluent and plausible text. However, these models are prone to produce factually incorrect or unverifiable statements, often called hallucinations. This impedes their deployment in real-world applications for information-seeking contexts Mallen et al. (2022). Prior work on hallucinations in natural language generation (NLG) often assumes the presence of a specific reference text and focuses on studying faithfulness to the references Ji et al. (2023). On the contrary, escalating apprehensions have been articulated concerning LM generations that are not grounded in any specific source text, but rather in world knowledge and facts Zhang et al. (2023).
Figure 1: Overview of taxonomy for LM hallucinations and our model, Fava, a fine-grained detection system.
@@
Several recent work studies automatic hallucination detection Min et al. (2023) or editing outputs Gao et al. (2022) to address such LM hallucinations. These systems typically categorize hallucinations into simplistic binary distinctions like factual or not factual (Figure 1 top). We argue that hallucinations manifest in diverse forms, each requiring varying degrees of careful assessments to verify factuality. Entity-level contradictions are usually evident and can be easily rectified with a single reference. Conversely, errors involving fabricated entities (e.g., The Messi Diaries in Figure 1) demand thorough verification across multiple sources. This underscores the need for a more fine-grained approach to detect hallucination for model development and human verification.
In this work, we propose automatic fine-grained hallucination detection (Figure 1 bottom), a new task requiring a system to provide precise identification of hallucination sequences, discern different types of hallucinations based on a taxonomy, and suggest refinements. We focus on hallucinations in information-seeking scenarios, where grounding to world knowledge matters. Our taxonomy (Figure 2) hierarchically classifies hallucinations in LM generations into six categories, based on careful pilot studies with NLP experts.
We construct a new fine-grained hallucination benchmark, FavaBench, by carefully annotating approximately 1,000 responses of three widely used LMs (Llama2-Chat 7B, 70B and ChatGPT [^2] We use gpt-3.5-turbo-0301 throughout this work.) to diverse knowledge-intensive queries. Each response is annotated at the span level to identify hallucinations, including erroneous subspace, types, and potential refinements. Our analysis reveals all models include at least one hallucination in the majority of their responses (e.g., 70.2% in Llama2 7B and 59.8% in ChatGPT). In addition to the widely-studied entity-level errors, other error types like unverifiable sentences make up over 60% of LM-generated hallucinations, highlighting the urgent need to study and detect such diverse types.
We introduce Fava, a new retrieval-augmented LM that can identify and mark hallucinations at the span level using a unified syntax (Figure 1, left bottom). Due to the annotation costs, we design an LM-based synthetic data generation process, and train Fava on the 35k resulting instances. We compare Fava with state-of-the-art LMs on fine-grained hallucination detection and editing in LM outputs. Fava significantly outperforms ChatGPT with and without external knowledge by 23.7% on fine-grained hallucination detection. Furthermore, Fava effectively corrects hallucinations in diverse LM outputs—Fava editing improves FActScore Min et al. (2023) of Alapaca 7, 13B, and ChatGPT by 4.4, 9.3 and 3.3%, respectively.
@@ 2 Related Work
Type Example ChatGPT Llama2-7b Llama2-70b
@@
 Relation Lionel Messi acquired was acquired by Paris Saint-Germain. 4.7% 3.4% 2.8%
  Sentence Messi has yet to be captain for the Argentina football team. 18.9% 15.7% 19.9%
 Invented Messi is known for his famous airplane kick technique. 14.2% 17.4% 9.5%
@@
 Unverifiable In his free time, Messi enjoys singing songs for his family. 9.6% 6.3% 6.3%
Table 1: Distribution of different errors across ChatGPT, Llama2-Chat-7B and Llama2-Chat-70B outputs and examples of hallucinations for each type from our human-annotated data
Hallucinations in NLG.
Prior taxonomies proposed for summarization Pagnoni et al. (2021) or text simplifications Devaraj et al. (2022) often assume the existence of specific source text and focus on faithfulness to the source Ji et al. (2023). As this is qualitatively different from LMs’ factual hallucinations grounded in world knowledge, we extend prior work and build a new taxonomy for hallucinations in LM outputs.
Detecting and editing hallucinations in LMs.
Several recent or concurrent studies propose methods to predict factuality by giving binary labels of a statement being factual or not Manakul et al. (2023); Min et al. (2023), or focusing on entity-level hallucination editing Gao et al. (2022); Chen et al. (2023). Recent surveys have attempted to categorize hallucinations in LM generations, often stretching the definition of hallucinations to broad error types in LM-generated text, which cannot be traced to specific spans in text (Huang et al., 2023; Zhang et al., 2023; Rawte et al., 2023). We propose a fine-grained taxonomy for hallucinations in long-form text generation for information-seeking context to improve factual consistency. While prior work often develops binary factuality verification systems on top of proprietary LMs Chern et al. (2023), we train a fine-grained detection system using our carefully designed data generation pipeline.
Fact verification for human-written claims.
While research on hallucination detection focuses on identifying errors in model-generated text, a related area of research focuses on identifying factual inaccuracies in human written claims (Bekoulis et al., 2021). Many large-scale datasets have been proposed for fact verification given Wikipedia Thorne et al. (2018), scientific articles (Wadden et al., 2020) or news documents (Wang, 2017). Several systems have been developed for fact-checking in these settings (Thorne and Vlachos, 2018; Schuster et al., 2021; Nakov et al., 2021), but have not been tested for LM-generated text. While our work focuses on hallucination detection in LM generations, Fava can be adapted to fact-check human written claims.
@@ 3.1 Hallucination Taxonomy
Based on our definition of hallucinations, we build a hierarchical taxonomy to categorize them into fine-grained error types. Inspired by prior task-specific taxonomies (Pagnoni et al., 2021; Devaraj et al., 2022), we introduce new categories to describe more complex errors surfacing in LM generations. We conducted a pilot annotation with nine NLP experts to ensure good coverage across diverse error types. Figure 2 shows our taxonomies to classify LM hallucinations and Table 1 shows examples of each type of hallucination. We group hallucinations into two major categories: statements that contradict world knowledge (Type 1) and unverifiable statements (Types 2, 3, and 4).
(1a)  Entity : Contradictory entity errors are a sub-category within contradictory statement errors where an entity in a statement is incorrect and changing that single entity can make the entire sentence factually correct.
@@
(1c)   Sentence : Contradictory sentence errors refer to cases where a full statement entirely contradicts relevant evidence from the web.
@@
(3)  Subjective : Subjective errors refer to an expression that lacks universal validity and is often influenced by personal beliefs, opinions, or biases.
(4)  Unverifiable : This refers to statements where the LM output contains facts, but none of the retrieved evidence from the web can directly support or contradict the fact (e.g., personal matters). [^4] While  Invented and  Unverifiable are highly related,  Invented are the hallucinations where we can verify that some core entities or subjects of the sentences don’t exist.
@@ 3.2 Tasks and Metrics
Figure 3: Distribution of hallucination types in ChatGPT, Llama2-Chat-7B and Llama2-Chat-70B outputs across four datasets of diverse information-seeking queries based on our human annotated benchmark.
Task 1: Fine-grained hallucination detection.
In this task, the system is expected to identify fine-grained errors in an LM output. Due to the subjectivity of span-level annotations, for automatic evaluation, we evaluate systems’ abilities to detect whether an error type $t$ exists in a sentence $s_{i}\in y$. Given an output $y$ consisting of $L$ sentences, we assume the availability of ground-truth error type annotations $e^{*t}_{i}\in\{\texttt{TRUE},\texttt{FALSE}\}$, which is a binary label of an error type $t$ existing in the $i$th sentence (TRUE) or not (FALSE). Following the fact verification literature Thorne et al. (2018); Schuster et al. (2021); Feng et al. (2023) of computing precision and recall, we measure those metrics while we extend them for each error type. For each type, a system predicts $e^{t}_{i}$ and we evaluate precision and recall as follows:
$\text{Prec}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=\texttt{TRUE}]}$ (1)
$\text{Recall}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{*t}_{i}=\texttt{TRUE}]}$ (2)
@@
Fine-grained detection can be simplified into a binary classification task, in which a system predicts if a sentence $s_{i}$ includes any factual errors or not. [^5] This is similar to FActScore Min et al. (2023), while we predict and evaluate factuality at the sentence level without extracting atomic facts.
Task 2: Hallucination editing.
Some hallucinations can be fixed with minimal span-level edits, while others need to be removed or marked as unverifiable. In the editing task, the system is expected to suggest fine-grained edits to improve the factuality of the LM output. We evaluate a system’s ability to improve the factuality of given output $y$, by comparing scores estimated by off-the-shelf systems, such as FActScore Min et al. (2023), as: $f(\hat{y})-f(y)$, where $f$ indicates an estimated factuality scores and $\hat{y}$ indicates edited output.
@@ 4 Benchmark: FavaBench
We create FavaBench to facilitate model development and evaluation in fine-grained hallucination detection and understand how prevalent those different types of hallucinations are. FavaBench consists of multi-way fine-grained annotations on three LM responses to queries in multiple domains.
Source prompts.
The source prompts include a collection of 200 information-seeking queries, spanning four different data sources. See examples in Appendix Table 4.
  • Knowledge-intensive queries sampled from the Open Assistant dataset Köpf et al. (2023), by prompting GPT-4 to judge whether each query from the dataset requires world knowledge.
  • A sample of Open QA prompts from the No Robots dataset Rajani et al. (2023).
  • Instruction-following prompts requiring more reasoning and knowledge curated by the authors.
  • Synthetically-created prompts 50 prompts that require more fine-grained knowledge, by converting data-to-text WebNLG dataset Gardent et al. (2017) using a template.
Annotation details and quality.
We obtain responses to the collected source prompts using ChatGPT (gpt-3.5-turbo-0301;  Ouyang et al. 2022) Llama2-Chat 7B and Llama2-Chat 70B Touvron et al. (2023) in a zero-shot manner and collect 600 responses to our diverse information-seeking prompts. We recruited 20 students (ten undergraduate and ten NLP graduate students) to annotate the factual accuracy of the responses based on our proposed taxonomy. Each instance is annotated by two annotators who completed 45-minute in-person and virtual training sessions. [^6] Annotating each response takes approximately 10 minutes, and we pay USD 3.5 for each annotation. Our annotation interface and details are in Appendix A. To validate our annotation quality, we calculated inter-annotator agreement using Cohen kappa scores and found that our annotators had high agreement in detection across passages, with 75.1% agreement in detection at the sentence level and 60.3% agreement in exact error type detection at the sentence level across passages. Appendix Table 6 shows examples of our annotation results.
Analysis on annotated data.
Figure 3 presents a detailed breakdown of distributions across fine-grained categories in the three domains. 59.8%, 70.2% and 64.9% of the responses of ChatGPT, Llama2-Chat 7B and Llama2-Chat 70B include at least one hallucination, respectively.  Entity  is the most widely recognized error type, making up 48.3% of the detected errors. Yet, there are diverse types of errors prevalent like invented statements or contradictory statements, which make up 14.1% and 18.1% of the detected errors, respectively. The error distributions vary across different source prompts and LMs.  Invented  are more common in Llama2-Chat models. Fewer errors in No Robots subsets than other subsets may be because their seed prompts are less knowledge-intensive, or ask about popular factual knowledge (e.g., How long was the Revolutionary War?), which is often memorized by LMs Mallen et al. (2022).

### v4

@@ Abstract
Large language models (LMs) are prone to generate factual errors, which are often called hallucinations. In this paper, we introduce a comprehensive taxonomy of hallucinations and argue that hallucinations manifest in diverse forms, each requiring varying degrees of careful assessments to verify factuality. We propose a novel task of automatic fine-grained hallucination detection and construct a new evaluation benchmark, FavaBench, that includes about one thousand fine-grained human judgments on three LM outputs across various domains. Our analysis reveals that ChatGPT and Llama2-Chat (70B, 7B) exhibit diverse types of hallucinations in the majority of their outputs in information-seeking scenarios, highlighting the need to build fine-grained systems. To this end, we train Fava, a powerful retrieval-augmented LM by carefully creating synthetic data to detect and correct fine-grained hallucinations. Our automatic and human evaluations show that Fava significantly outperforms retrieval-augmented ChatGPT and GPT-4 on fine-grained hallucination detection. Furthermore, Fava outperforms widely-used hallucination detection systems on binary detection and shows effectiveness in editing to improve the factuality. [^1] Our code, data, and demo are available at https://fine-grained-hallucination.github.io/.
Figure 1: Overview of our taxonomy, fine-grained hallucination detection task, and Fava.
@@ 1 Introduction
Large language models (LMs; Brown et al. 2020) can generate highly fluent and plausible text. However, these models are prone to produce factually incorrect or unverifiable statements, often called hallucinations. This impedes their deployment in real-world applications for information-seeking contexts (Mallen et al., 2022; Asai et al., 2024). Prior work on hallucinations in natural language generation (NLG) often assumes the presence of a specific reference text and focuses on studying faithfulness to the references (Ji et al., 2023). On the contrary, escalating apprehensions have been articulated about LM generations that are not grounded in any specific source text, but rather in world knowledge (Zhang et al., 2023).
@@
Several recent work studies automatic hallucination detection (Min et al., 2023) or editing outputs (Gao et al., 2022) to address such LM hallucinations. These systems typically categorize hallucinations into simplistic binary distinctions like factual or not factual (Figure 1 right) or mainly focus on entity-level errors. We argue that hallucinations manifest in diverse forms, each requiring varying degrees of careful assessments to verify factuality. Entity-level contradictions are usually evident and can be easily rectified with a single reference. Conversely, errors involving fabricated entities (e.g., The Messi Diaries in Figure 1) demand thorough verification across multiple sources. This underscores the need for a more fine-grained approach to detect hallucinations.
In this work, we propose automatic fine-grained hallucination detection (Figure 1 left), a new task requiring a system to provide precise identification of hallucination sequences, discern different types based on a taxonomy, and suggest refinements. We focus on hallucinations in information-seeking scenarios, where grounding to world knowledge matters. Our taxonomy (Figure 2) hierarchically classifies hallucinations in LM generations into six categories. Our taxonomy is based on iterative pilot studies with NLP experts, to ensure comprehensive coverage of all significant types of hallucinations.
We construct a new fine-grained hallucination benchmark, FavaBench, by carefully annotating approximately 1,000 responses of three widely used LMs (Llama2-Chat 7B, 70B and ChatGPT [^2] We use gpt-3.5-turbo-0301 throughout this work.) to diverse knowledge-intensive queries. Each response is annotated at the span level to identify hallucinations, including erroneous subspace, types, and potential refinements. Our analysis reveals all models include at least one hallucination in the majority of their responses (e.g., 70.2% in Llama2 7B and 59.8% in ChatGPT). Besides the widely-studied entity-level errors, other types like unverifiable sentences make up over 60% of LM-generated hallucinations, highlighting the urgent need for fine-grained detection.
Yet, prior methods often only provide statement-level binary hallucination predictions. To advance fine-grained hallucination detection, we introduce Fava, a new retrieval-augmented LM that can identify and mark hallucinations at span-level using a unified syntax (Figure 1, left). We design an LM-based synthetic data generation and train Fava on the 35k resulting instances. We compare Fava with state-of-the-art LMs on fine-grained hallucination detection based on FavaBench. Fava significantly outperforms ChatGPT with and without external knowledge by 23.7% on fine-grained hallucination detection. On binary hallucination detection, Fava outperforms a widely-used system, FActScore (Min et al., 2023) in addition to other strong baselines leveraging GPT4 or ChatGPT. Fava also effectively corrects hallucinations in diverse LM outputs—Fava editing improves a factuality score  (Min et al., 2023) of Alapaca 7, 13B, and ChatGPT by 4.4, 9.3 and 3.3%, respectively.
@@ 2 Related Work
Type Example Chat L-7b L-70b
@@
 Relation Lionel Messi acquired was acquired by PSG. 4.7% 3.4% 2.8%
  Sentence Messi was never captain for Argentina football team. 18.9% 15.7% 19.9%
 Invented Messi is known for his famous airplane kick. 14.2% 17.4% 9.5%
@@
 Unverifiable When free, Messi enjoys singing songs for his family. 9.6% 6.3% 6.3%
Table 1: Distribution of different errors across ChatGPT (Chat), Llama2-Chat-7B (L-7B) and Llama2-Chat-70B (L-70B) outputs and examples of hallucinations for each type.
Hallucinations in NLG. Prior taxonomies proposed for summarization (Pagnoni et al., 2021), text simplifications (Devaraj et al., 2022) or knowledge-grounded dialogue (Dziri et al., 2022) often assume the existence of specific source text and focus on faithfulness to the source (Ji et al., 2023). As this is qualitatively different from LMs’ factual hallucinations grounded in world knowledge, we extend prior work and build a new taxonomy for hallucinations in LM outputs.
Detecting and editing hallucinations in LMs. Several recent or concurrent studies propose methods to predict factuality by giving binary labels of a statement being factual or not (Manakul et al., 2023; Min et al., 2023), focusing on entity-level hallucination editing (Gao et al., 2022; Chen et al., 2023), or concentrating purely on detection (Li et al., 2024). Recent surveys have attempted to categorize hallucinations in LM generations, often stretching the definition of hallucinations to broad error types in LM-generated text, which cannot be traced to specific spans in text (Huang et al., 2023; Zhang et al., 2023; Rawte et al., 2023). We propose a fine-grained taxonomy for hallucinations in long-form text generation for information-seeking context to improve factual consistency. While prior work often develops binary factuality verification systems on top of proprietary LMs (Chern et al., 2023), we train a fine-grained detection system using our carefully designed data generation pipeline.
Fact verification for human-written claims. While research on hallucination detection focuses on identifying errors in model-generated text, a related area of research focuses on identifying factual inaccuracies in human written claims (Bekoulis et al., 2021). Many large-scale datasets have been proposed for fact verification given Wikipedia (Thorne et al., 2018), scientific articles (Wadden et al., 2020) or news documents (Wang, 2017). Several systems have been developed for fact-checking in these settings (Thorne & Vlachos, 2018; Schuster et al., 2021; Nakov et al., 2021), but have not been tested for LM-generated text. While our work focuses on hallucination detection in LM generations, Fava can be adapted to fact-check human written claims.
@@ 3.1 Hallucination Taxonomy
Based on our definition of hallucinations, we build a hierarchical taxonomy to categorize them into fine-grained error types. Inspired by prior task-specific taxonomies (Pagnoni et al., 2021; Devaraj et al., 2022), we introduce new categories to describe more complex errors surfacing in LM generations. We conducted a pilot annotation with nine NLP experts to ensure good coverage across diverse error types. Figure 2 shows our taxonomies to classify LM hallucinations. Table 1 shows examples of each type of hallucination. We group hallucinations into two major categories: statements that contradict world knowledge (Type 1) and unverifiable statements (Types 2, 3, and 4).
(1a)  Entity : Contradictory entity errors are a sub-category within Type 1, where an entity in a statement is incorrect and changing that single entity can make the entire sentence factually correct.
@@
(1c)   Sentence : Contradictory sentence errors refer to cases where a full statement entirely contradicts relevant evidence from the web, and cannot be solved via phrase-level edits.
@@
(3)  Subjective : Subjective errors refer to expressions about existing entities that lack universal validity. These statements often do not contain facts and are influenced by personal beliefs or opinions.
(4)  Unverifiable : These are statements where the LM output contains facts, but no retrieved evidence from the web can directly support or contradict the fact (e.g., private details).  [^4] Differing from  Subjective errors which focus on opinionated statements lacking facts,  Unverifiable specifically focuses on fact-based statements that just lack evidence.  [^5] While  Invented and  Unverifiable are highly related,  Invented are the hallucinations where we can verify that some core entities or subjects of the sentences don’t exist.
@@ 3.2 Tasks and Metrics
Figure 3: Distribution of hallucination types in ChatGPT, Llama2-Chat-7B and Llama2-Chat-70B outputs across four datasets of diverse information-seeking queries.
Task 1: Fine-grained hallucination detection. In this task, the system is expected to identify fine-grained errors in an LM output. Due to the subjectivity of span-level annotations, for automatic evaluation, we evaluate systems’ abilities to detect whether an error type $t$ exists in a sentence $s_{i}\in y$. Given an output $y$ consisting of $L$ sentences, we assume the availability of ground-truth error type annotations $e^{*t}_{i}\in\{\texttt{TRUE},\texttt{FALSE}\}$, which is a binary label of an error type $t$ existing in the $i$th sentence (TRUE) or not (FALSE). Following the fact verification literature (Thorne et al., 2018; Schuster et al., 2021; Feng et al., 2023) of computing precision and recall, we measure those metrics while we extend them for each error type. For each type, a system predicts $e^{t}_{i}$ and we evaluate precision and recall as:
$\text{Prec}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=\texttt{TRUE}]}{\rm,~{}~{}~{}}\text{Recall}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{*t}_{i}=\texttt{TRUE}]}$ (1)
@@
Fine-grained detection can be simplified into a binary classification task, in which a system predicts if a sentence $s_{i}$ includes any factual errors or not. [^6] This is similar to FActScore (Min et al., 2023), while we predict and evaluate factuality at the sentence level without extracting atomic facts. We compare Fava against FActScore on this task.
Task 2: Hallucination editing. Some hallucinations can be fixed with minimal span-level edits, while others need to be removed or marked as unverifiable. In the editing task, the system is expected to suggest fine-grained edits to improve the factuality of the LM output. We evaluate a system’s ability to improve the factuality of given output $y$, by comparing scores estimated by off-the-shelf systems, such as FActScore (Min et al., 2023), as: $f(\hat{y})-f(y)$, where $f$ indicates an estimated factuality scores and $\hat{y}$ indicates edited output.
@@ 4 Benchmark: FavaBench
FavaBench consists of around 1k fine-grained annotations on three LM responses to queries in multiple domains.
Source prompts. The source prompts include a collection of 200 information-seeking queries, spanning four different data sources. See examples in Appendix Table 6.
  • 50 Knowledge-intensive queries from Open Assistant (Köpf et al., 2023), by prompting GPT-4 to judge whether each query from the dataset requires world knowledge.
  • 50 Open QA prompts from the No Robots dataset (Rajani et al., 2023).
  • 50 instructions requiring more reasoning and knowledge by the authors.
  • 50 synthetically-created prompts that require more fine-grained knowledge, by converting data-to-text WebNLG dataset (Gardent et al., 2017) using a template.
Annotation details and quality. We obtain responses to the collected source prompts using ChatGPT (gpt-3.5-turbo-0301;  Ouyang et al. 2022) Llama2-Chat 7B and Llama2-Chat 70B (Touvron et al., 2023) in a zero-shot manner and collect 600 responses to our diverse information-seeking prompts. We recruited 20 students (ten undergraduate and ten NLP graduate students) to annotate the factual accuracy of the responses based on our proposed taxonomy. Each instance is annotated by two annotators who completed 45-minute in-person and virtual training sessions. [^7] Annotating each response takes approximately 10 minutes, and we pay USD 3.5 for annotation. We provide annotators with detailed instructions, examples, and training to ensure high-quality annotations. Our annotation interface and details are in Appendix A and examples in Appendix Table 8. To validate our annotation quality, we calculated inter-annotator agreement using Cohen kappa scores. Our annotators have high agreement in detection across passages, with 75.1% agreement in detection at the sentence level and 60.3% agreement in exact error type detection.
Analysis on annotated data. Figure 3 presents a detailed breakdown of distributions across fine-grained categories in the three domains. 59.8%, 70.2% and 64.9% of the responses of ChatGPT, Llama2-Chat 7B and Llama2-Chat 70B include at least one hallucination, respectively.  Entity  is the most widely recognized error type, making up 48.3% of the detected errors. Yet, there are diverse types of errors prevalent like invented statements or contradictory statements, which make up 14.1% and 18.1% of the detected errors, respectively. The error distributions vary across different source prompts and LMs.  Invented  are more common in Llama2-Chat models. Fewer errors in No Robots subsets than other subsets may be because their seed prompts are less knowledge-intensive, or ask about popular factual knowledge (e.g., How long was the Revolutionary War?), which is often memorized by LMs (Mallen et al., 2022).